# Cross-Problem Solving for Network Optimization: Is Problem-Aware Learning the Key?

### v1

@@ IV-C Problem-Aware Diffusion
We design a Transformer-based LDM [27] as the neural network component of PAD, with the Transformer [28] architecture as its backbone. Specifically, LDM is characterized by its ability to project variable-dimensional inputs from the data space into a fixed-dimensional latent space via an encoder, and then perform diffusion-based noise adding and denoising in this latent space. As illustrated in Fig. 2, our element-wise encoder maps each individual element in the input vector $\mathbf{x}$ and solution vector $\mathbf{y}$ into a token vector, in a way analogous to how a text encoder tokenizes characters. To account for the distinct physical meanings of different elements in $\mathbf{x}$ and $\mathbf{y}$, we assign an integer label to each element representing its physical type (e.g., power, channel gain), allowing the encoder to generate discriminative embeddings. Let the encoder and decoder networks be parameterized by $\boldsymbol{\theta}_{\mathrm{E}}$ and $\boldsymbol{\theta}_{\mathrm{D}}$, respectively. The training objective of the encoder-decoder is given as $f_{\boldsymbol{\theta}_{\mathrm{D}}}(f_{\boldsymbol{\theta}_{\mathrm{E}}}(\mathbf{x}))=\mathbf{x}$, where we apply mean squared error (MSE) loss on the element values and cross-entropy loss on the associated physical type labels. The training procedure is detailed in Algorithm 1.
Input: $N$ training problems and their corresponding datasets $\boldsymbol{\mathcal{D}}=\{\mathcal{D}_{1},\ldots,\mathcal{D}_{i},\ldots,\mathcal{D}_{N}\}$ and problem embedding after ranked pooling $\{\mathbf{e}^{\mathrm{p}}_{1},\ldots,\mathbf{e}^{\mathrm{p}}_{i},\ldots,\mathbf{e}^{\mathrm{p}}_{N}\}$, the constraint-aware module $\boldsymbol{\theta}_{\mathrm{C}}$, constraint judgment tool function $\mathrm{if\_satisfy\_cons}()$;
@@
5                   $\mathbf{x}$leftarrow batch data from $\mathcal{D}_{i}$;
6                   $\mathbf{y}$leftarrow randomly sampled solutions;
7                   $\tilde{\mathbf{r}}$leftarrow $f_{\boldsymbol{\theta}_{\mathrm{C}}}(\mathbf{e}^{\mathrm{p}}_{i},\mathbf{x},\mathbf{y})$;
8                   $\mathbf{r}$leftarrow $\mathrm{if\_satisfy\_cons}(\mathbf{e}^{\mathrm{p}}_{i},\mathbf{x},\mathbf{y})$;
9                   $\mathcal{L}$leftarrow $\mathrm{CrossEntropyLoss}(\mathbf{r},\tilde{\mathbf{r}})$;
@@ 10                   Backpropagation on $\boldsymbol{\theta}_{\mathrm{C}}$ with $\mathcal{L}$;
For the backbone of PAD, we incorporate the ranked pool mechanism introduced in Section IV-B. As shown in Fig. 3, the encoded sequence of $\mathbf{x}$ is concatenated with the pooled problem embedding sequence obtained via ranked pool, forming a condition sequence $\mathbf{c}$ that guides the denoising process.
During training, the encoded optimal solution sequence $\mathbf{z}_{0}$ is progressively noised according to a diffusion schedule up to time step $T$, eventually yielding a pure Gaussian noise vector $\mathbf{z}_{T}$. The Transformer-based diffusion backbone $\boldsymbol{\theta}_{\mathrm{B}}$ takes the noisy vector $\mathbf{z}_{t}$ and the condition sequence $\mathbf{c}$ as input, and predicts the noise to be removed. The model is trained using MSE loss between the predicted and actual noise.
Here, $\mathbf{c}$ and $\mathbf{z}_{t}$ correspond to the source and target sequences for the Transformer, respectively. At inference time, we sample $\mathbf{z}_{T}$ directly from Gaussian noise and iteratively apply $\boldsymbol{\theta}_{\mathrm{B}}$ to denoise it back to $\tilde{\mathbf{z}_{0}}$ without noise, which is then decoded to obtain the final generated solution $\tilde{\mathbf{y}}$.
IV-D Constraint-Aware Module
Input: $N$ training problems and their corresponding datasets $\boldsymbol{\mathcal{D}}=\{\mathcal{D}_{1},\ldots,\mathcal{D}_{i},\ldots,\mathcal{D}_{N}\}$ and problem embedding after ranked pooling $\{\mathbf{e}^{\mathrm{p}}_{1},\ldots,\mathbf{e}^{\mathrm{p}}_{i},\ldots,\mathbf{e}^{\mathrm{p}}_{N}\}$, diffusion step $T$, constraint loss start epoch $e$, the trained models $\boldsymbol{\theta}_{\mathrm{C}}$, $\boldsymbol{\theta}_{\mathrm{E}}$ and $\boldsymbol{\theta}_{\mathrm{D}}$;
@@ IV-D Constraint-Aware Module
5                   $\mathbf{x},\mathbf{y}^{*}$leftarrow batch data from $\mathcal{D}_{i}$;
6                   $\mathbf{c}_{\mathrm{x}},\mathbf{z}_{0}$leftarrow $f_{\boldsymbol{\theta}_{\mathrm{E}}}(\mathbf{x},\mathbf{y}^{*})$;
7                   $\mathbf{c}$leftarrow $\mathrm{cat}(\mathbf{e}^{\mathrm{p}}_{i},\mathbf{c}_{\mathrm{x}})$;
8                   $\mathbf{t}$leftarrow sample time step from $\mathrm{Uniform}(1,T)$;
9                   $\boldsymbol{\epsilon}$leftarrow sample noise from $\mathcal{N}(0,\mathbf{I})$;
10                   $\mathbf{z}_{t}$leftarrow noise adding on $\mathbf{z}_{0}$ based on $\boldsymbol{\epsilon}$ and $\mathbf{t}$;
11                   $\tilde{\boldsymbol{\epsilon}}$leftarrow $f_{\boldsymbol{\theta}_{\mathrm{B}}}(\mathbf{c},\mathbf{t},\mathbf{z}_{t})$;
12                   $\mathcal{L}_{\mathrm{eps}}$leftarrow $\|\boldsymbol{\epsilon}-\tilde{\boldsymbol{\epsilon}}\|^{2}$;
13                   $\mathcal{L}_{\mathrm{cons}}$leftarrow 00;
14                   if $\mathrm{in\ the\ last}\ e\ \mathrm{epochs}$ then
15                         $\mathbf{z}_{t},\tilde{\boldsymbol{\epsilon}}$leftarrow select $\mathbf{z}_{t}$ and the corresponding $\tilde{\boldsymbol{\epsilon}}$ where $t<\frac{T}{2}$;
16                         $\tilde{\mathbf{z}_{0}}$leftarrow denoising on $\mathbf{z}_{t}$ using $\tilde{\boldsymbol{\epsilon}}$;
17                         $\mathcal{L}_{\mathrm{cons}}$leftarrow $f_{\boldsymbol{\theta}_{\mathrm{C}}}(\mathbf{e}^{\mathrm{p}}_{i},\mathbf{x},f_{\boldsymbol{\theta}_{\mathrm{D}}}(\tilde{\mathbf{z}_{0}}))$;
18
19                   end if
20                  $\mathcal{L}$leftarrow $\mathcal{L}_{\mathrm{eps}}+\mathcal{L}_{\mathrm{cons}}$;
@@ 21                   Backpropagation on $\boldsymbol{\theta}_{\mathrm{B}}$ with $\mathcal{L}$;
To improve the likelihood of generating feasible solutions, we design an independent constraint-aware module based on a TransformerEncoder [20] and a classification head. This module functions as a binary classifier that takes the input vector $\mathbf{x}$, the solution $\mathbf{y}$, and the pooled problem embedding as input, and predicts whether $\mathbf{y}$ violates the hard constraints of the corresponding problem (outputting $1$ if it does, and 00 otherwise). By approximating non-differentiable constraint functions through a differentiable neural network, this module introduces constraint violation penalties during the solution generation process, guiding the model toward producing valid solutions. This also increases the diversity and quality of feasible solutions discovered during parallel sampling. Augmenting the GDMs’ noise prediction loss with additional preference-based objectives has proven effective in the AI community [48]. The training process of the constraint-aware module is detailed in Algorithm 2.
@@
In implementation, we prepend a classification (CLS) token [49] to the input sequence and use its output representation for downstream classification. As provided in Algorithm 3, to mitigate the potential gradient interference between the noise prediction loss $\mathcal{L}_{\mathrm{eps}}$ and the constraint violation loss $\mathcal{L}_{\mathrm{cons}}$, we apply $\mathcal{L}_{\mathrm{cons}}$ only during the final phase of training. By introducing $\mathcal{L}_{\mathrm{cons}}$ after $\mathcal{L}_{\mathrm{eps}}$ has largely converged, we enable more efficient optimization of $\mathcal{L}_{\mathrm{cons}}$ while avoiding the convergence instability that may arise from joint training from the beginning. Moreover, we compute $\mathcal{L}_{\mathrm{cons}}$ only for the final solutions derived from the denoising steps in the latter half of the diffusion process. This strategy ensures that the incorporation of constraints does not excessively disturb the denoising trajectory.
@@ V-A Experimental Settings
Figure 4: The similarity score matrices between problems calculated on problem embeddings extracted using different pooling methods, where scores of two problems close to 1 represent high similarity and scores close to 0 represent low similarity.
@@ V-A1 Datasets
We construct three variants (with 5, 6, and 7 optimization variables) for each of the six problems introduced in Section III-B. For each of these 18 problem settings, we produce a training set of 60,000 samples and a test set of 2,000 samples using the Gurobi solver [50]. All samples in the dataset are optimal solutions, and the detailed ranges of input parameters are provided in our open-source repository.
For clarity in presenting evaluation results, we refer to the six problems in Section III-B by short names: [“P1”, …, “P6”]. Combined with their optimization dimensions, the test sets are named accordingly. For example, “P1_D5” denotes the test set for problem P1 with 5-dimensional optimization variables. We also assign each of the 18 problem settings a unique problem index ranging from 1 to 18. For instance, problem index 1 corresponds to “P1_D5”, and index 4 to “P2_D5”.
The original problem feature embeddings are generated using MathBERT [25], with each problem represented in LaTeX mathematical format as input. The resulting embeddings typically range from 100 to 200 tokens in length, with each token having a dimensionality of 768. Although the token sequence lengths of DeepSeekMath [26] are comparable to those of MathBERT, the dimensionality per token reaches 4096 due to differences in its tokenizer or vocabulary. This leads to a prohibitively high computational cost in the token compression layer after pooling, and thus DeepSeekMath is omitted from subsequent experiments. Unless otherwise specified, we adopt our proposed ranked pool method with $m=20$ for pooling the problem embeddings in the following experiments.
To analyze the model’s cross-problem generalization ability, we partition the six problems from Section III-B into training and testing subsets using two distinct strategies. Both partitioning strategies are based on inter-problem similarity scores derived from pooled problem embeddings, as described in Section V-B. Following the approach in Section IV-B, we compute the total similarity score of each problem with all others and sort them accordingly. In the first strategy, referred to as “ascending”, we select problems with the lowest total similarity scores for training. To ensure at least one unseen problem is left for evaluation, we construct five different training settings under this strategy. In the second strategy, referred to as “descending”, we instead select problems with the highest total similarity scores for training, resulting in five different training settings. Therefore, we explore 10 different training configurations for the backbone based on these two problem selection strategies.
All datasets were generated on a machine with two Intel(R) Xeon(R) Silver 4410Y processors. The problem was that embedding generation and model inference were both performed on an RTX 4090 24 GB.

### v2

@@ IV-C Problem-Aware Diffusion
We design a Transformer-based LDM [28] as the neural network component of PAD, with the Transformer [29] architecture as its backbone. Specifically, LDM is characterized by its ability to project variable-dimensional inputs from the data space into a fixed-dimensional latent space via an encoder, and then perform diffusion-based noise adding and denoising in this latent space. As illustrated in Fig. 2, our element-wise encoder maps each individual element in the input vector $\mathbf{x}$ and solution vector $\mathbf{y}$ into a token vector, in a way analogous to how a text encoder tokenizes characters. To account for the distinct physical meanings of different elements in $\mathbf{x}$ and $\mathbf{y}$, we assign an integer label to each element representing its physical type (e.g., power, channel gain), allowing the encoder to generate discriminative embeddings. These physical types are mainly designed to enable the element-wise encoder–decoder to produce parameter and variable embeddings with physical meaning, thereby allowing the latent diffusion model to learn specialized, high-quality solution distributions for data with different physical interpretations. In practice, however, the latent diffusion model is not required to achieve perfect classification of variable types; instead, it focuses on learning the distribution of high-quality variable values, which avoids the excessive difficulty of simultaneously generating both variable values and types.
Let the encoder and decoder networks be parameterized by $\boldsymbol{\theta}_{\mathrm{E}}$ and $\boldsymbol{\theta}_{\mathrm{D}}$, respectively. The training objective of the encoder-decoder is given as $f_{\boldsymbol{\theta}_{\mathrm{D}}}(f_{\boldsymbol{\theta}_{\mathrm{E}}}(\mathbf{x}))=\mathbf{x}$, where we apply mean squared error (MSE) loss on the element values and cross-entropy loss on the associated physical type labels. The training procedure is detailed in Algorithm 1.
Input: NN training problems and their corresponding datasets $\boldsymbol{\mathcal{D}}=\{\mathcal{D}_{1},\ldots,\mathcal{D}_{i},\ldots,\mathcal{D}_{N}\}$ and complete problem embedding after rank pooling $\{\mathcal{E}^{\mathrm{p,eq}}_{1},\ldots,\mathcal{E}^{\mathrm{p,eq}}_{i},\ldots,\mathcal{E}^{\mathrm{p,eq}}_{N}\}$, the constraint-aware module $\boldsymbol{\theta}_{\mathrm{C}}$, constraint judgment tool function $\mathrm{if\_satisfy\_cons}()$;
@@
5        $\mathbf{x}$ $\leftarrow$ batch data from $\mathcal{D}_{i}$;
6        $\mathbf{y}$ $\leftarrow$ randomly sampled solutions;
7        $\mathbf{e}_{i}$ $\leftarrow$ randomly selected from $\mathcal{E}^{\mathrm{p,eq}}_{i}$;
8        $\tilde{\mathbf{r}} \leftarrow$ $f_{\boldsymbol{\theta}_{\mathrm{C}}}(\mathbf{e}_{i},\mathbf{x},\mathbf{y})$;
9        $\mathbf{r}$ $\leftarrow$ $\mathrm{if\_satisfy\_cons}(\mathbf{e}_{i},\mathbf{x},\mathbf{y})$;
[…]
@@ 11        Backpropagation on $\boldsymbol{\theta}_{\mathrm{C}}$ with $\mathcal{L}$;
For the backbone of PAD, under the $i$-th problem, we incorporate the rank pool mechanism introduced in Section IV-B. As shown in Fig. 3, the encoded sequence of $\mathbf{x}_{i}$ is concatenated with the pooled problem embedding sequence obtained via rank pool, forming a condition sequence $\mathbf{c}$ that guides the denoising process.
During training, the encoded optimal solution sequence $\mathbf{z}^{0}$ is progressively noised according to a diffusion schedule up to time step $T$, eventually yielding a pure Gaussian noise vector $\mathbf{z}^{T}$. The Transformer-based diffusion backbone $\boldsymbol{\theta}_{\mathrm{B}}$ takes the noisy vector $\mathbf{z}^{t}$ and the condition sequence $\mathbf{c}$ as input, and predicts the noise to be removed. The model is trained using MSE loss between the predicted and actual noise.
Here, $\mathbf{c}$ and $\mathbf{z}^{t}$ correspond to the source and target sequences for the Transformer, respectively. At inference time, we sample $\mathbf{z}^{T}$ directly from Gaussian noise and iteratively apply $\boldsymbol{\theta}_{\mathrm{B}}$ to denoise it back to $\tilde{\mathbf{z}^{0}}$ without noise, which is then decoded to obtain the final generated solution $\tilde{\mathbf{y}_{i}}$.
Input: NN training problems and their corresponding datasets $\boldsymbol{\mathcal{D}}=\{\mathcal{D}_{1},\ldots,\mathcal{D}_{i},\ldots,\mathcal{D}_{N}\}$ and complete problem embedding after rank pooling $\{\mathcal{E}^{\mathrm{p,eq}}_{1},\ldots,\mathcal{E}^{\mathrm{p,eq}}_{i},\ldots,\mathcal{E}^{\mathrm{p,eq}}_{N}\}$, constraint missing problem embedding after rank pooling $\{\mathcal{E}^{\mathrm{p,miss}}_{1},\ldots,\mathcal{E}^{\mathrm{p,miss}}_{i},\ldots,\mathcal{E}^{\mathrm{p,miss}}_{N}\}$, diffusion step $T$, constraint loss start epoch $e$, the trained models $\boldsymbol{\theta}_{\mathrm{C}}$, $\boldsymbol{\theta}_{\mathrm{E}}$ and $\boldsymbol{\theta}_{\mathrm{D}}$;
@@
5        $\mathbf{x},\mathbf{y}^{*}$ $\leftarrow$ batch data from $\mathcal{D}_{i}$;
6        $\mathbf{c}_{\mathrm{x}},\mathbf{z}_{0}$ $\leftarrow$ $f_{\boldsymbol{\theta}_{\mathrm{E}}}(\mathbf{x},\mathbf{y}^{*})$;
7        $\mathbf{e}_{i}$ $\leftarrow$ randomly selected from $\mathcal{E}^{\mathrm{p,eq}}_{i}\bigcup\mathcal{E}^{\mathrm{p,miss}}_{i}$;
8        $\mathbf{c}$ $\leftarrow$ $\mathrm{cat}(\mathbf{e}_{i},\mathbf{c}_{\mathrm{x}})$;
9        $\mathbf{t}$ $\leftarrow$ sample time step from $\mathrm{Uniform}(1,T)$;
10        $\boldsymbol{\epsilon}$ $\leftarrow$ sample noise from $\mathcal{N}(0,\mathbf{I})$;
11        $\mathbf{z}_{t}$ $\leftarrow$ noise adding on $\mathbf{z}_{0}$ based on $\boldsymbol{\epsilon}$ and $\mathbf{t}$;
12        $\tilde{\boldsymbol{\epsilon}} \leftarrow$ $f_{\boldsymbol{\theta}_{\mathrm{B}}}(\mathbf{c},\mathbf{t},\mathbf{z}_{t})$;
13        $\mathcal{L}_{\mathrm{eps}}$ $\leftarrow$ $\|\boldsymbol{\epsilon}-\tilde{\boldsymbol{\epsilon}}\|^{2}$;
14        $\mathcal{L}_{\mathrm{cons}}$ $\leftarrow$ 0;
15        if in the last $e$ epochs and $\mathbf{e}_{i}$ from $\mathcal{E}^{\mathrm{p,eq}}_{i}$ then
16           $\mathbf{z}_{t},\tilde{\boldsymbol{\epsilon}} \leftarrow$ select $\mathbf{z}_{t}$ and the corresponding $\tilde{\boldsymbol{\epsilon}}$ where $t<\frac{T}{2}$;
17           $\tilde{\mathbf{z}_{0}} \leftarrow$ denoising on $\mathbf{z}_{t}$ using $\tilde{\boldsymbol{\epsilon}}$;
18           $\mathcal{L}_{\mathrm{cons}}$ $\leftarrow$ $)f_{\boldsymbol{\theta}_{\mathrm{C}}}(\mathbf{e}_{i},\mathbf{x},f_{\boldsymbol{\theta}_{\mathrm{D}}}(\tilde{\mathbf{z}_{0}}$));
19
20       $\mathcal{L}$ $\leftarrow$ $\mathcal{L}_{\mathrm{eps}}+\mathcal{L}_{\mathrm{cons}}$;
@@ IV-D Constraint-Aware Module
To improve the likelihood of generating feasible solutions, we design an independent constraint-aware module based on a TransformerEncoder [21] and a classification head, which is used only during training. This module functions as a binary classifier that takes the parameter $\mathbf{x}$, the solution $\mathbf{y}$, and the pooled problem embedding as input, and predicts whether $\mathbf{y}$ violates the hard constraints of the corresponding problem (outputting $1$ if it does, and 0 otherwise). By approximating non-differentiable constraint functions through a differentiable neural network, this module introduces constraint violation penalties during the solution generation process, guiding the model toward producing valid solutions. This also increases the diversity of feasible solutions discovered during parallel sampling. Augmenting the GDMs’ noise prediction loss with additional preference-based objectives has proven effective in the AI community [56]. The training process of the constraint-aware module is detailed in Algorithm 2.
@@
In implementation, we prepend a classification (CLS) token [57] to the input sequence and use its output representation for downstream classification. To mitigate the potential gradient interference between the noise prediction loss $\mathcal{L}_{\mathrm{eps}}$ and the constraint violation loss $\mathcal{L}_{\mathrm{cons}}$, we apply $\mathcal{L}_{\mathrm{cons}}$ only during the final phase of training. By introducing $\mathcal{L}_{\mathrm{cons}}$ after $\mathcal{L}_{\mathrm{eps}}$ has largely converged, we enable more efficient optimization of $\mathcal{L}_{\mathrm{cons}}$ while avoiding the convergence instability that may arise from joint training from the beginning. Moreover, we compute $\mathcal{L}_{\mathrm{cons}}$ only for the final solutions derived from the denoising steps in the latter half of the diffusion process. This strategy ensures that the incorporation of constraints does not excessively disturb the denoising trajectory.
To enhance the model’s sensitivity to constraints and robustness to the diversity of problem embeddings, we design a constraint-gated loss training mechanism inspired by the idea of classifier-free guidance (CFG) [58]. This mechanism prevents the constraint signal from being treated merely as background information of input parameters and thus overlooked by the model. Specifically, for each problem ($i$), we construct two sets of embeddings during training, denoted as $\mathcal{E}^{\mathrm{p,eq}}_{i}$ and $\mathcal{E}^{\mathrm{p,miss}}_{i}$. All embeddings in $\mathcal{E}^{\mathrm{p,eq}}_{i}$ are equivalent to the original problem, differing only in the ordering of terms in the objective and constraint functions or in variable signs. In contrast, the embeddings in $\mathcal{E}^{\mathrm{p,miss}}_{i}$ not only involve such permutations and sign variations, but also include incomplete constraint specifications, where parts or even all of the constraints may be missing. Both $\mathcal{E}^{\mathrm{p,eq}}_{i}$ and $\mathcal{E}^{\mathrm{p,miss}}_{i}$ contain 3–5 different embeddings each, which are used exclusively during training. At each iteration, the model randomly samples an embedding from their union, and the constraint-aware module performs forward computation of the constraint loss only when the input comes from $\mathcal{E}^{\mathrm{p,eq}}_{i}$. This also reduces the computational overhead of automatic differentiation and backpropagation through the constraint-aware module. It is worth noting that our constraint-gated loss differs fundamentally from CFG. The constraint-gated loss is determined solely by the content of the diffusion model’s input condition rather than by the presence or absence of a problem embedding. Our design is primarily intended as an exploration of a novel training strategy to strengthen the model’s awareness of constraints, rather than to reproduce the expected effects of CFG’s dual noise-prediction and weighted-sum mechanism. During the inference phase, embeddings are introduced only from $\mathcal{E}^{\mathrm{p,eq}}_{i}$.
@@ V-A Experimental Settings
TABLE I: Default settings for PAD’s main hyperparameters.
Model Parameter Value
encode_dims [4, 32, 128]
hidden_dim 128
decode_dims [128, 32]
type_num 19
encoder_decoder_train_epochs 20
encoder_decoder_init_lr $1.5\times 10^{-4}$
transformer_encoder_num 6
transformer_decoder_num 6
attention_header_num 8
transformer_dropout 0.1
constraint_module_train_epochs 30
constraint_module_init_lr $1.5\times 10^{-4}$
pool_type rank pool
pool_length ($m$) 20
diffusion_type Gaussian
diffusion_steps 50
inference_diffusion_steps 5
parallel_sampling_num ($p_{\rm num}$) 2
constrain_loss_start_ratio 0.5
use_gated_constraint_loss True
diffusion_training_epochs 70
diffusion_init_lr $2\times 10^{-4}$
@@ V-A1 Datasets
We label the ten problems in Section III-B sequentially as [P1, …, P10]. For each problem, we set three output variable dimensions and generate optimal samples using the Gurobi solver [59]. Except for P7, where the three dimensions are set to 3, 4, and 5 (corresponding to total output dimensions of 6, 8, and 10) due to the excessive runtime of the solver on higher-dimensional inputs, all other problems use dimensions of 5, 6, and 7 (corresponding to total output dimensions of 10, 12, and 14 in the mixed-integer case). The values of input parameters are configured with reference to realistic settings from related literature, as detailed in our open-source repository.
The $\mathcal{E}^{\mathrm{p,eq}}_{i}$ and $\mathcal{E}^{\mathrm{p,miss}}_{i}$ are obtained by first embedding each LaTeX-format expression with MathBERT [26] to produce raw problem embeddings, and then applying pooling operations. The resulting embeddings typically range from 100 to 200 tokens in length, with each token having a dimensionality of 768. Although the token sequence lengths of DeepSeekMath [27] are comparable to those of MathBERT, the dimensionality per token reaches 4096 due to differences in its tokenizer or vocabulary. This leads to a prohibitively high computational cost in the token compression layer after pooling, and thus DeepSeekMath is omitted from subsequent experiments. Unless otherwise specified, we adopt our proposed rank pool method with $m=20$ for pooling the problem embeddings in the following experiments.
To analyze the cross-problem generalization ability of the model, we follow the clustering in Section III-B and select one problem from each cluster (namely P1, P4, and P7) together with the independent P10 to form the complete training dataset. Each problem contributes 60,000 samples per dimension. Since the difficulty variations within each cluster are not significant, choosing P1, P4, and P7 as representatives is an empirical decision, but it is sufficient to control for variables and support the intended experiments. For evaluation, we use 2,000 non-overlapping test samples per dimension for every problem.
All datasets were generated on a machine equipped with an Intel(R) i7-13700F processor and 16 GB of RAM. The problem embedding generation, model training, and inference were all performed on an RTX 4090 24 GB.